OpenStudy (anonymous):

E is the sum of the first 40 positive even integers.K is the sum of the first 40 positive odd integers. What is E-K?

OpenStudy (anonymous):

40

OpenStudy (anonymous):

40 becoz for each pair of one even and one odd no the difference is 1 so for 40 pairs it is 40

OpenStudy (nowhereman):

Indeed, or in formula \[E - K = \sum_{n=1}^{40}{(2n+1)} - \sum_{n=1}^{40}{2n} = \sum_{n=1}^{40}{1} = 40\]

OpenStudy (anonymous):

first even no is 2 and first odd no is 1 so difference is 1 second even no is 4 and second odd no is 3, again difference is 1 so for 40 pairs, it is 40

OpenStudy (anonymous):

\[E=\sum_{n=1}^{n}2n=n(n+1)/2\] \[K=\sum_{1}^{n}(2n-1)=n(n+1)-n\] so E-K=n and this holds for any n

OpenStudy (anonymous):

here u have n=40 so ans is 40